1. Let Upper T Subscript Upper A Baseline colon Upper R Superscript 3 Baseline right-arrow Upper R Superscript 3 be multiplication by A and find the dimension of the subspace of Upper R Superscript 3 consisting of all vectors x for which Upper T Subscript Upper A Baseline left-parenthesis x right-parenthesis equals 0.

a. A= 1 1 0
1 0 3
1 0 3
The dimension is????

b. A = 1 4 0
1 4 0
1 4 0
The dimension is???

c. A = 3 0 0
-3 3 0
3 3 3
The dimension is????
1
Expert's answer
2020-01-10T13:25:04-0500

"Ax=0, \quad x= \begin{bmatrix}
x_1
\\ x_2
\\ x_3
\end{bmatrix}"

"a) A = \begin{bmatrix}
1&1&0
\\ 1&0&3
\\ 1&0&3
\end{bmatrix}"

"Ax=0 \; \Leftrightarrow \begin{cases} x_1+x_2=0 \\ x_1+3x_3=0 \\ x_1+3x_3=0 \end{cases}"

"x_1=-x_2=-3x_3"

"\begin{bmatrix}
x_1
\\ x_2
\\ x_3
\end{bmatrix}
= \begin{bmatrix}
-3x_3
\\ 3x_3
\\ x_3
\end{bmatrix}
= x_3\begin{bmatrix}
-3
\\ 3
\\ 1
\end{bmatrix}, \forall x_3 \in \mathbb{R}"

"\Rightarrow dimension \; of \; sub space \; of \; \mathbb{R}^3 \; consisting \; of \; all \; vectors \; x \; for \; which \; Ax=0"

"equals \; 1."

"b) A = \begin{bmatrix}
1&4&0
\\ 1&4&0
\\ 1&4&0
\end{bmatrix}"

"Ax=0 \; \Leftrightarrow \begin{cases} x_1+4x_2=0 \\ x_1+4x_2=0 \\ x_1+4x_2=0 \end{cases}"

"x_1=-4x_2"

"\begin{bmatrix}
x_1
\\ x_2
\\ x_3
\end{bmatrix}
= \begin{bmatrix}
-4x_2
\\ x_2
\\ x_3
\end{bmatrix}
= x_2\begin{bmatrix}
-4
\\ 1
\\ 0
\end{bmatrix}+x_3 \begin{bmatrix}
0
\\ 0
\\ 1 \end{bmatrix} \forall x_2,x_3 \in \mathbb{R}"

"\Rightarrow dimension \; of \; sub space \; of \; \mathbb{R}^3 \; consisting \; of \; all \; vectors \; x \; for \; which \; Ax=0"

"equals \; 2."

"c) A = \begin{bmatrix}
3&0&0
\\ -3&3&0
\\ 3&3&3
\end{bmatrix}"

"Ax=0 \; \Leftrightarrow \begin{cases} x_1=0 \\ -3x_1+3x_2=0 \\ 3x_1+3x_2+3x_3=0 \end{cases}"

"x_2=x_1=0, x_3=-x_1-x_2=0"

"\begin{bmatrix}
x_1
\\ x_2
\\ x_3
\end{bmatrix} = \begin{bmatrix}
0
\\ 0
\\ 0
\end{bmatrix}"

"\Rightarrow dimension \; of \; sub space \; of \; \mathbb{R}^3 \; consisting \; of \; all \; vectors \; x \; for \; which \; Ax=0"

"equals \; 0."